this is a question on my math homework. determine if the following lines are parallel, perpendicular, or neither. explain how your know show any work.
first line
-2x=18-4x i figured the top line out -2y=4x=18 divided all by -2 y=-2x-9 slope is -2.

second line--im stumped on this one.
1y/5-2x/5=1/20

Answer by Alan3354(69443)

determine if the following lines are parallel, perpendicular, or neither. explain how your know show any work.
first line
-2x=18-4x i figured the top line out -2y=4x=18 divided all by -2 y=-2x-9 slope is -2.

second line--im stumped on this one.
1y/5-2x/5=1/20
Multiply by 5
y - 2x = 1/4
y = 2x + 1/4
Slope = +2
---------------
Not parallel, not perpendicular